Child pages
  • Engineering Report Day 12 TasMars2013
Skip to end of metadata
Go to start of metadata

Engineering Report: 02/06/2013
Emma Braegen

 
Was Power Interrupted:
no
Time Of Power Outage:

Length Of Power Outage:

Diesel Fuel Level:
0.17
Propane Fuel Level:
0.68
Gasoline Fuel Level:
0.5
Fuel Notes And Comments:

Atv Oil Quarts:
N.A.
Car Oil Quarts:
N.A.
Oil Notes And Comments:

Trailer Water Fill Needed:
yes
Trailer Water Refilled:
no
Potable Water Meter Reading:
62732.3
Water Notes And Comments:

Greenhouse Notes And Comments:

Telescope Used:
no
Observatory Notes And Comments:

H A Lp Bugs:

Network Failed:
no
Network Throttled:
no
Network Signal Strength:
187
Network Download Allowance:
29
Network Notes And Comments:
Internet bandwidth seems to be doing well today, still not sure what the usage problem was previously
Hvac Air Conditioner Used:
no
Hvac Furnace Used:
yes
Hvac Lab Wall Heater Used:
no
Hvac Notes And Comments:

Hab Car Used:
no
Hab Car Oil Checked:
no
Hab Car Tire Status:

Hab Car Odometer:

Hab Car Notes And Comments:

Atv One Used:
no
Atv One Oil Checked:
no
Atv One Fuel Used Gals:

Atv One Tires Status:

Atv One Hours Used Day:

Atv One Hours Used Year:

Atv One Notes And Comments:

Atv Two Used:
no
Atv Two Oil Checked:
no
Atv Two Fuel Used Gals:

Atv Two Tires Status:

Atv Two Hours Used Day:

Atv Two Hours Used Year:

Atv Two Notes And Comments:

Atv Three Used:
no
Atv Three Oil Checked:
no
Atv Three Fuel Used Gals:

Atv Three Tires Status:

Atv Three Hours Used Day:

Atv Three Hours Used Year:

Atv Three Notes And Comments:

Atv Four Used:
no
Atv Four Oil Checked:
no
Atv Four Fuel Used Gals:

Atv Four Tires Status:

Atv Four Hours Used Day:

Atv Four Hours Used Year:

Atv Four Notes And Comments:

Atv Five Used:
no
Atv Five Oil Checked:
no
Atv Five Fuel Used Gals:

Atv Five Tires Status:

Atv Five Hours Used Day:

Atv Five Hours Used Year:

Atv Five Notes And Comments:

Comments:

Only engineering issue today is working with space suit number 2 is difficult since one of the tube is broken at one point. It is taped together but occasionally this comes apart and needs retaping. The rear latch on the helment is also broken. These are only minor issues however and the suit is still usable.

 
I did another test for the dust suit experiment today with Melanie using one of the space suits and trying to collect as much dust and dirt as possible onto painters overalls. Turns out it is not as easy to do as I originally thought but hopefully I can get some useful results by comparing the dirt and dust brought in from EVA with the material found around the hab.
The trailer water tank is mostly empty now, I could probably get a few more centimeters of water out of it but the external tank is full at the moment.

  • No labels